If g on the surface of the earth is 9.8m/s2. Find its value at depth of 3200km

Grade:12th pass

1 Answers

Arun
6 years ago
Dear,g=GM/r^2 , let the volume mass density of earth be =p,g=G4/3pi.r^3.1/r^2=G.4pi.r/3When radius is halved g`=G.2pi.r/3=g/2Hence, g=4.9m/s^2 at depth of 3200 km